# ArcMemo: Abstract Reasoning Composition with Lifelong LLM Memory

### v2

@@ Appendix D Additional Related Work
Test-time learning (sometimes referred to as online learning) describes methods that updates predictions at inference time Bottou & Cun (2003); Bottou & Le Cun (2005). However, trends towards (1) massive model sizes and (2) black box models accessed through APIs make continually updating parameters computationally impractical or downright impossible. In light of this, LLM test-time learning research looked to pursue parameter-free adaptation by making use of LLMs’ remarkable capacity for instruction following and self-reflection. Methods such as Reflexion (Shinn et al., 2023), Self-Refine (Madaan et al., 2023), and Self-RAG (Asai et al., 2023), inter alia attempt to correct mistakes at test time by way of self-reflection. Follow-up work (Huang et al., 2024; Kamoi et al., 2024) find that self-correction methods depend on the availability of some extrinsic feedback or verification mechanism to reliably improve end performance and that these methods produce dubious results absent these grounded sources of feedback. Still, self-reflection remains a key component of self-evolving agents (Gao et al., 2025).
@@
Memory also plays a crucial role in LLM-based agents. Park et al. (2023) structured memory as a stream of episodic observations and distilled high-level summaries from them to guide decision-making. Contextual Experience Replay discretizes trajectories into “experiences” (containing environment dynamics from a trajectory and skills storing actions from said trajectory) and retrieves relevant blocks to guide new episodes (Liu et al., 2025). HiAgent organizes working memory (within a single query) hierarchically by subgoals, summarizing and replacing traces to improve long-horizon efficiency (Hu et al., 2024). External-memory systems such as Mem0 and MemP manage a persistent store with explicit add–update–prune operations (MemP casts control as an MDP), yielding gains on dialogue and tool-use/planning tasks (Chhikara et al., 2025; Fang et al., 2025). MemP in particular is a concurrent work that also demonstrates the efficacy of abstraction, although they target rewriting action sequences as step sequences in explicitly agentic settings, contrasting with ArcMemo’s general problem solving focus. Voyager uses LLMs to autonomously acquire and store skills while exploring Minecraft, with a growing library of reusable code snippets written by the agent itself Wang et al. (2023a). Other systems, like the ResearchAgent, use memory initialized from domain-specific corpora, such as scientific papers, to ground query responses in high-utility prior knowledge (Baek et al., 2024).
Figure 4: Token Efficiency Plot: Comparing various settings’ official score (oracle@2) on the public validation subset to the tokens used by the reasoning model for strictly puzzle solving

### v3

@@ Appendix E Additional Related Work
Test-time learning (sometimes referred to as online learning) describes methods that update predictions at inference time Bottou & LeCun (2003; 2005). However, trends towards (1) massive model sizes and (2) black box models accessed through APIs make continually updating parameters computationally impractical or downright impossible. In light of this, LLM test-time learning research sought to pursue parameter-free adaptation by using LLMs’ remarkable capacity for instruction following and self-reflection. Methods such as Reflexion (Shinn et al., 2023), Self-Refine (Madaan et al., 2023), and Self-RAG (Asai et al., 2023), inter alia attempt to correct mistakes at test-time by way of self-reflection. Follow-up work (Huang et al., 2024; Kamoi et al., 2024) finds that self-correction methods depend on the availability of some extrinsic feedback or verification mechanism to reliably improve end performance and that these methods produce dubious results absent these grounded sources of feedback. Still, self-reflection remains a key component of self-evolving agents (Gao et al., 2025).
@@
Memory also plays a crucial role in LLM-based agents. Park et al. (2023) structured memory as a stream of episodic observations and distilled high-level summaries from them to guide decision-making. Contextual Experience Replay discretizes trajectories into “experiences” (containing environment dynamics from a trajectory and skills storing actions from said trajectory) and retrieves relevant blocks to guide new episodes (Liu et al., 2025). HiAgent organizes working memory (within a single query) hierarchically by subgoals, summarizing and replacing traces to improve long-horizon efficiency (Hu et al., 2024). External-memory systems such as Mem0 and MemP manage a persistent store with explicit add–update–prune operations (MemP casts control as an MDP), yielding gains on dialogue and tool-use/planning tasks (Chhikara et al., 2025; Fang et al., 2025). MemP in particular is a concurrent work that also demonstrates the efficacy of abstraction, although it targets rewriting action sequences as step sequences in explicitly agentic settings, contrasting with ArcMemo’s general problem-solving focus. Voyager uses LLMs to autonomously acquire and store skills while exploring Minecraft, with a growing library of reusable code snippets written by the agent itself Wang et al. (2023a). Other systems, like the ResearchAgent, use memory initialized from domain-specific corpora, such as scientific papers, to ground query responses in high-utility prior knowledge (Baek et al., 2024).